OpenStudy (anonymous):

solve for g. p=3g+3v

OpenStudy (texaschic101):

p = 3g + 3v --- subtract 3v p - 3v = 3g --- divide by 3 (p - 3v)/3 = g or p/3 - v = g

OpenStudy (anonymous):

i got both of those wrong

OpenStudy (texaschic101):

I do not know why....it looks right to me

OpenStudy (texaschic101):

@Mertsj ...whats up with this question

OpenStudy (mertsj):

The answer is (p-3v)/3

OpenStudy (texaschic101):

thats the right answer ....I do not know how she got it wrong

OpenStudy (mertsj):

The asker should double check and make absolutely certain that it is posted correctly. If it is the asker should review the rules for entry of answers

OpenStudy (texaschic101):

thanks for taking a look martsj :)

OpenStudy (mertsj):

The parentheses are absolutely essential.

OpenStudy (mertsj):

yw

OpenStudy (texaschic101):

oops..typo...mertsj...sorry bout that